A projectile is launched with V0 = 8.3 m/s and initial angle = 0.83 radians above the horizontal. What is the initial horizontal component of the projectile velocity in miles per hour?
Answer #1

The velocity has 2 components.

Horizontal velocity component Vx= v* cos(theta)

Vx = 8.3 * Cos(0.83)

Vx=5.6m/s

Now m/s into mile/hr

1m/s=2.237mph

So 5.6m/s= 5.6*2.237=12.52mph